tele deah being given beloo ·ghe yields ar givenin c (43) A)A Raced on ha ihes the ayent an
Answer #1

Answer:

Crop replication after analysing the high yield variety crop becomes essential for producing the high productivity crops .

For this , different methods are employed such as critical difference in 1 block of data set and analysing the data at various blocks and then choosing the high yield crops .

In this case we have 4 replication / blocks -

In 1st block - Avg. yield ( A , B , C ) = 48+47+49/3 = 48

In 2nd block - Avg. yield ( A , B, C ) = 51+49+53/3 = 51

In 3rd block - Avg . yield ( A , B, C ) = 52+52+49 /3 = 51

In 4th block - Avg. yield ( A,B,C) = 49+51+50/3 = 50

We can see that in 2nd and 3rd replication avg. Yield is highest , so these sets of crop variety are suitable for replication .

There is not any significant difference in the avg . Yield of the crops as the difference between the highest and lowest average yield is only about 3 kg .

The layout of the 2nd and 3rd variety crop are suitable for the crop replication and will yield better results in crop productivity.

Please rate the answer good if you like it.